OpenStudy (anonymous):

3. Find the following product: (x - 3)(x + 3) I think that I have found the answer to this question but can someone give me a second opinion to basically help my decision.... I am kind of stuck between two answers on it... PLEASE help me

OpenStudy (dumbcow):

Distribute or FOIL --> x^2 + 3x -3x -9 = x^2 -9

OpenStudy (anonymous):

I get confused on the FOIL method... can u tell me how that goes... my teacher confuses me big time with that method...

OpenStudy (anonymous):

More in depth: F- multiply first terms \[X x X = X^2\] O - multipy outside terms \[X \times 3 = 3x\] I - Inside terms \[X x -3 = -3x\] L - Last terms \[3 x -3 = -9\] Put that all together and you get \[X^2 -3x + 3x -9\] when you combine like terms, that's \[X^2 -9\]

OpenStudy (anonymous):

sorry, some of the x's are supposed to be multiplication signs, but i'm sure you can figure which.

OpenStudy (anonymous):

lol... yea i figured that part out.... thank you for explaining that to me more in depth bc i was kind of lost there on that.... However, I had already put x^2 - 9 on my test before I asked for help or another opinion so I am happy that I got the question correct... LOL :)

OpenStudy (anonymous):

If needed, use the box method (example in attached file) (sorry for the bad handwriting, it's done on with a mouse.)

OpenStudy (anonymous):

Thank you so much... never heard of that before but will take note of that definitely...